# Fast, robust and laser-free universal entangling gates for trapped-ion quantum computing

### v2

@@ Abstract
A novel two-qubit entangling gate for trapped-ion quantum processors is proposed theoretically and demonstrated experimentally. During the gate, double-dressed quantum states are created by applying a phase-modulated continuous driving field. The speed of this gate is an order of magnitude higher than that of previously demonstrated radio frequency (RF) controlled two-qubit entangling gates in static magnetic field gradients. At the same time, the field driving the gate, dynamically decouples the qubits from amplitude and frequency noise, increasing the qubits’ coherence time by three orders of magnitude. The gate requires only a single continuous RF field per qubit, making it well suited for scaling a quantum processor to large numbers of qubits. Implementing this entangling gate, we generate the Bell states |$\ket{\Phi^{+}}$ and |$\ket{\Psi^{+}}$ in $\leq 313$ $\mathrm{\mathrm{\mu}}$s with fidelities up to $98^{+2}_{-3}\leavevmode\nobreak$in a static magnetic gradient of only 19.09 T/m. At higher magnetic field gradients, the entangling gate speed can be further improved to match that of laser-based counterparts.
@@ I Introduction
Trapped atomic ions are among physical platforms well suited for quantum information processing [1, 2]. Outstanding performance has been realized in trapped-ion quantum processors by using focused lasers to coherently control individual ionic qubits [3, 4, 5, 6, 7, 8]. Trapped ions coherently controlled by radio frequency (RF) signals are particularly suited for scalability, since technological challenges associated with using laser light for coherent control of qubits are avoided in this laser-free approach [9, 10, 11, 12, 13]. High fidelity single- and two-qubit gates [14, 15, 16, 17, 18], as well as complete quantum algorithms [19, 20] have been realized with RF-controlled ions. Single-qubit rotations are performed with low cross-talk [21], which is crucial for fault-tolerant quantum computing.
@@
Conditional quantum dynamics with trapped ion qubits using RF radiation requires a magnetic field gradient to couple internal qubit states to the ions’ motional states [9, 10, 22]. RF-based coherent control can be achieved either by far-field RF radiation applied in a local static magnetic field gradient [13, 15, 19], or by a dynamic magnetic field gradient [12, 4, 17]. While the two approaches share important features, for instance, laser-free coherent manipulation and magnetic field gradient induced coupling (MAGIC), each approach offers distinct advantages when it comes to scaling up ion-based quantum computing.
@@
In both RF-based approaches, two-qubit gate speed has lagged behind its laser-based counterparts [17, 15]. Here, we introduce and experimentally realize a novel two-qubit entangling gate with gate speed an order of magnitude faster than previous RF-based gates in static magnetic field gradients [15] and a factor of two faster than previous two-qubit gates using the dynamic gradient approach [17]. Furthermore, in contrast to previous work, only a single phase modulated RF field per ion is required for its operation, further simplifying the scaling up of trapped-ion quantum computers.
@@
RF-controlled qubits in static magnetic gradients often have a relatively short coherence time due to magnetic field fluctuations. Common practice to counteract decoherence is dynamical decoupling [23, 24]. Inspired by nuclear magnetic resonance refocusing pulses [25, 26, 27], refocusing techniques, referred to as pulsed dynamical decoupling, were designed for quantum registers to increase coherence time [28, 29, 30, 31, 32, 33, 34, 35]. The ideas of dynamical decoupling have also been extended to protect various gate operations by composite pulses [36, 37, 38, 39, 40]. Alternatively, dynamical decoupling is achieved by applying a continuous driving field, rather than pulses. Continuous dynamical decoupling, too, has been successfully incorporated into entangling gates in trapped ions [41, 42, 43, 15]. The gate presented in this work utilizes continuous dynamical decoupling to create and entangle double-dressed qubit states [44, 45], which offer built-in robustness against external noise making additional dynamical decoupling superfluous. Using these double-dressed qubits, we improved the coherence time by three orders-of-magnitude compared to the bare qubits.
@@
The gate implemented here can be interpreted as a geometric phase gate [46, 47, 48, 49, 50, 51] between double-dressed qubits. A variant of double-dressed two-qubit gates has been theoretically proposed for spin qubits associated with silicon vacancy centers in diamond [52]. To our knowledge, the work presented here is the first experimental demonstration of a two-qubit gate based on double-dressed states in any physical platform.
@@
This article is structured as follows: After a brief description of the experimental apparatus in Sec. II, we give an account of the experimental investigation of double dressing individually addressed qubits in a two-ion crystal in Sec. III. Then, the theoretical description of the entangling gate driven by a phase modulated field is presented in Sec. IV.1, before we describe the experimental procedure and results for this two-qubit entangling gate in Sec. IV.2 and IV.3, respectively. We conclude this article with a discussion of phase modulated vs. amplitude modulated driving fields for entangling gates as well as a summary of results and an outlook on future work in Sec. V.
@@ II Experimental overview
We investigate double dressed state dynamics of a single qubit and implement conditional quantum dynamics with two qubits using 171Yb+ ions trapped in a macroscopic linear Paul trap with radial and axial center-of-mass (COM) angular frequencies of $2\pi\times 380$ kHz and $\nu=2\pi\times 98.08$ kHz, respectively. The experimental setup is similar to [18] and only briefly described here. The two qubits are encoded in hyperfine states of the electronic ground state of 171Yb+ ions, $|0\rangle\equiv{|^{2}\rm S_{1/2},\rm F=0,\rm m_{\rm F}=0\rangle}$ and $|1\rangle\equiv{|^{2}\rm S_{1/2},\rm F=1,\rm m_{\rm F}=-1\rangle}$. A static magnetic field gradient of $19.09(1)$ T/m applied along the trap axis (z-axis) yields individual Zeeman-shifts of the two qubits and results in individual angular resonance frequencies, $\omega_{0}^{(1)}$ and $\omega_{0}^{(2)}$, of the magnetic dipole transition between qubit states. Both frequencies are near $2\pi\times 12.6$ GHz and differ by $2\pi\times 4.342(1)$ MHz. Thus, qubits are individually coherently controlled with low cross-talk using global RF radiation [21].
The entangling gate proposed and implemented here takes advantage of a state selective force induced by the static magnetic gradient field coupling the internal qubit states to the axial vibrational states of the 2-ion crystal [9]. The coupling strength between qubit states and motional states, the latter acting as a quantum bus between qubits [1], is given by the effective Lamb-Dicke parameter, $\eta=g_{F}\mu_{B}(\partial_{z}B)\nu^{-3/2}/\sqrt{2Nm_{\text{Yb}}}$ that is proportional to the magnetic field gradient, $\partial_{z}B$ [9], $g_{F}$ is the hyperfine Landé g-factor of ${}^{171}\text{Yb}^{+}$ (appendix A), $m_{\text{Yb}}$ is the mass of ${}^{171}\text{Yb}$, $\mu_{B}$ is Bohr’s magneton, and $N$ is the number of ions in a linear Coulomb crystal. In the experiments presented here $\eta=0.0329$. Therefore, when using magnetic gradient induced coupling (MAGIC), laser light is not required for implementing conditional gates with trapped ions.
The ions are cooled close to their motional ground state in two stages. Initial Doppler cooling is followed by RF sideband cooling of both vibrational modes giving a mean phonon number of 0.6(5) in the center-of-mass (COM) mode measured with sideband thermometry [53]. The heating rate of one trapped ion in the current experimental setup is 0.19(3) phonons/ms for this mode.
We use an Arbitrary Waveform Generator (AWG) to generate the phase-modulated RF driving fields that double dress both qubits [45, 54, 44].
@@ III Double dressing of single qubits
The MAGIC scheme in a static gradient field requires a magnetic sensitive qubit transition, making the qubits susceptible to decoherence due to magnetic field fluctuations. We demonstrate experimentally that the double dressing field acts as continuous dynamical decoupling and improves the coherence time significantly by suppressing the effect of addressing frequency fluctuations as well as amplitude fluctuations of the RF driving field. The idea of using double dressed states to counter noise has been established in previous works on single qubits in nitrogen-vacancy centers in diamond [55, 44, 56]. Here, we extend double dressing to individual trapped ions in a two-ion Coulomb crystal, and investigate the coupling of the qubit resonance to vibrational modes by the phase modulated driving field.
Figure 1: Illustration of the emergence of double-dressed states and the consequent protection against frequency and amplitude fluctuations. Thin double arrows indicate resonance frequencies. Thick double arrows indicate Rabi frequencies. The bare qubit is driven near resonance, $\omega_{0}$ with a time dependent detuning achieved by phase modulation according to Eq. (5) (left: Bare). In the dressed state basis (middle: Dressed), the drive with Rabi frequency $\Omega_{1}$ translates into the energy gap $\Omega_{1}$. Relative frequency fluctuations $\delta\omega_{0}$ between driving field and qubit are suppressed in the dressed basis for $\delta\omega_{0}\ll\Omega_{1}$. The phase modulation transforms into an effective on-resonance drive in the dressed basis, creating double dressed states (right: Double-Dressed). In the double-dressed frame, the effective second drive translates into the double-dressed states energy gap. Amplitude fluctuations in the first drive $\delta\Omega_{1}$ are suppressed in the double-dressed basis for $\delta\Omega_{1}\ll\Omega_{2}$.
The double dressed scheme is illustrated for one qubit in Fig. 1 and is described in appendix A in detail. In brief, the ionic qubit in a harmonic oscillator potential is modeled as [9]
@@
Here, $\sigma_{z}$ is the the Pauli operator describing the qubit’s eigenstates. We consider the center-of-mass (COM) vibrational mode of two ions and denote $b$ as the ladder operator of the motional quanta in the effective harmonic trapping potential. In the mathematical description in this article we set $\hbar=1$. In all experiments reported in this article, we trap two ions (i.e., $N=2$) while the number of ions simultaneously addressed by a dedicated RF driving field changes. For the single-qubit experiments discussed in this section as well as for the two-qubit experiments in section IV, undesired crosstalk between qubits is negligible [21].
We drive one ion near-resonance with a phase modulated field such that
@@
where $\Omega_{1}$ is the Rabi frequency of the RF drive, $\sigma_{x}$ a Pauli operator, and $\Omega_{2}$ is a parameter quantifying the phase modulation amplitude.
When the driving field is applied (Eq. (2)), we obtain the Hamiltonian,
$H_{I}=\frac{\Omega_{1}}{2}S_{z}+\frac{\Omega_{2}}{2}S_{x}\cos(\Omega_{1}t)+\nu b^{\dagger}b-\frac{\eta\nu}{2}S_{x}\left(b+b^{\dagger}\right),$ (3)
in the rotating frame w.r.t. $\omega_{0}\sigma_{z}/2+\Omega_{2}\sigma_{z}/2\cos\left(\Omega_{1}t\right)$. Here, $S_{x}=-\sigma_{z},S_{y}=\sigma_{y},S_{z}=\sigma_{x}$ are the dressed basis operators. Dressing suppresses fluctuations of the bare state energy gap $\delta\omega_{0}\ll\Omega_{1}$ caused by either frequency fluctuations of the driving field, or by fluctuations of magnetically sensitive bare qubit states which in turn may be caused by electric noise fields.
Phase modulation of the dressing field, in addition, suppresses fluctuations in the amplitude of $\Omega_{1}$: Phase modulation creates, in an appropriate rotating frame, an effective on-resonance second drive in the dressed basis. This second drive dresses the qubit a second time, and thus dynamically decouples amplitude noise fluctuations in the drive $\delta\Omega_{1}$, which would otherwise reduce the coherence time of the dressed qubit (Fig. 1). These fluctuations are suppressed as long as $\delta\Omega_{1}\ll\Omega_{2}$. Furthermore, undesired coupling of the qubit to the motional state, which leads to a decay in the Rabi oscillations, is suppressed as long as $\Omega_{2}\gg\eta\nu$ and $|\Omega_{1}-\nu|\ll\Omega_{2}\ll|\Omega_{1}+\nu|$ (appendix A gives a detailed description).
@@ III.1 Extending the coherence time
Figure 2: Relaxation time $T_{2,\rho}$ of a single qubit as a function of the modulation depth $\Omega_{2}/\Omega_{1}$. The qubit is initialized in state $|0\rangle$, which is a superposition state in the dressed state basis. The phase-modulated driving field is continuously applied and the coherence time $T_{2,\rho}$ is extracted from an exponential decay of the Rabi fringe contrast. At a modulation depth $\Omega_{2}/\Omega_{1}\approx 0.75$, the relaxation time is enhanced by three orders of magnitude compared to the scenario without second dressing field.
The impact of the phase modulated dressing field on a qubit’s coherence time is shown in Fig. 2. We study the relaxation of a single qubit in a two-qubit crystal initialized in state $|0\rangle$, while the phase-modulated driving field given in Eq. (2) is continuously applied. Note, that this is an equal super-position state in the dressed basis, therefore, by fitting the decay of the oscillations’ contrast to an exponential decay we can extract the coherence time, $T_{2,\rho}$ of the dressed state.
When no modulation is applied, we observe ordinary Rabi oscillations (not shown) with frequency $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, close to the trap frequency $\nu=2\pi\times 98.08\leavevmode\nobreak$kHz. At low modulation strength, the coherence time is mainly limited by the near-resonant coupling of the qubit to the COM motional mode. With increasing modulation amplitude, $\Omega_{2}/\Omega_{1}$, the coherence time of the dressed state, $T_{2,\rho}$ increases.
@@
For $\Omega_{2}/\Omega_{1}=0.75$, used for the entangling gate mechanism in this work, we observe three orders of magnitude improvement in $T_{2,\rho}$, up to 0.27(2) s. The slightly reduced coherence time for $\Omega_{2}/\Omega_{1}\approx 1$ is due to second-order fluctuations of $\delta\Omega_{2}$ [44] and to the quantum stark shift induced by counter rotating terms.
III.2 Coherent coupling to motional degrees-of-freedom

### v3

# Present address: ]eleQtron GmbH, Heeserstr. 5, 57072 Siegen,Germany

@@ Abstract
A novel two-qubit entangling gate for trapped-ion quantum processors is proposed theoretically and demonstrated experimentally. During the gate, double-dressed quantum states are created by applying a phase-modulated continuous driving field. The speed of this quantum gate is an order of magnitude higher than that of previously demonstrated rf controlled two-qubit entangling gates in static magnetic field gradients. At the same time, the field driving the gate dynamically decouples the qubits from amplitude and frequency noise, increasing the qubits’ coherence time by 3 orders of magnitude. The gate requires only a single continuous rf field per qubit, making it well suited for scaling a quantum processor to large numbers of qubits. Implementing this entangling gate, we generate the Bell states |$\ket{\Phi^{+}}$ and |$\ket{\Psi^{+}}$ in less than or equal to $313$ µs with fidelities up to $98^{+2}_{-3}$in a static magnetic gradient of only 19.09 T/m. At higher magnetic field gradients, the entangling gate speed can be further improved to match that of laser-based counterparts.
@@ I Introduction
Trapped atomic ions are among the physical platforms well suited for quantum information processing [1, 2]. Outstanding performance has been achieved in trapped-ion quantum processors using focused lasers to coherently control individual ionic qubits [3, 4, 5, 6, 7, 8]. Trapped ions coherently controlled by rf signals are particularly suitable for scalability, since the technological challenges associated with the use of laser light for coherent control of qubits are avoided in this laser-free approach [9, 10, 11, 12, 13]. High-fidelity single- and two-qubit gates [14, 15, 16, 17, 18], as well as complete quantum algorithms [19, 20], have been realized with rf-controlled ions. Single-qubit rotations are performed with low crosstalk [21], which is crucial for fault-tolerant quantum computing.
@@
Conditional quantum dynamics with trapped-ion qubits using rf radiation requires a magnetic field gradient to couple the internal qubit states to the ions’ motional states [9, 10, 22]. rf-based coherent control can be achieved either by far-field rf radiation applied in a local static magnetic field gradient [13, 15, 19] or by a dynamic magnetic field gradient [12, 4, 17]. Although the two approaches share important features—for instance, laser-free coherent manipulation and magnetic gradient induced coupling (MAGIC)—each approach offers distinct advantages when it comes to scaling up ion-based quantum computing.
@@
In both rf-based approaches, the two-qubit gate speed has lagged behind its laser-based counterparts [17, 15]. Here, we introduce and experimentally realize a novel two-qubit entangling gate with gate speed an order of magnitude faster than previous rf-based gates in static magnetic field gradients [15] and a factor of 2 faster than previous two-qubit gates using the dynamic gradient approach [17]. Furthermore, in contrast to previous work, only a single phase-modulated rf field per ion is required for its operation, further simplifying the scaling up of trapped-ion quantum computers.
@@
The rf-controlled qubits in static magnetic gradients often have a relatively short coherence time because of magnetic field fluctuations. Common practice to counteract decoherence is dynamical decoupling [23, 24]. Inspired by nuclear magnetic resonance refocusing pulses [25, 26, 27], refocusing techniques, referred to as pulsed dynamical decoupling, were designed for quantum registers to increase coherence time [28, 29, 30, 31, 32, 33, 34, 35]. The ideas of dynamical decoupling have also been extended to protect various gate operations by composite pulses [36, 37, 38, 39, 40]. Alternatively, dynamical decoupling is achieved by applying a continuous driving field, rather than pulses. Continuous dynamical decoupling has also been successfully incorporated into entangling gates in trapped ions [41, 42, 43, 15]. The gate presented in this work utilizes continuous dynamical decoupling to create and entangle double-dressed qubit states [44, 45], which offer built-in robustness against external noise, making additional dynamical decoupling superfluous. Using these double-dressed qubits, we improve the coherence time by 3 orders of magnitude compared to the bare qubits.
@@
The gate implemented here can be interpreted as a geometric phase gate [46, 47, 48, 49, 50, 51] between double-dressed qubits. A variant of double-dressed two-qubit gates has been theoretically proposed for spin qubits associated with silicon vacancy centers in diamond [52]. To the best of our knowledge, the work presented here is the first experimental demonstration of a two-qubit gate based on double-dressed states in any physical platform.
@@
This article is structured as follows. We provide a brief description of the experimental apparatus in Sec. II, we give an account of the experimental investigation of double dressing individually addressed qubits in a two-ion crystal in Sec. III. Then, the theoretical description of the entangling gate driven by a phase-modulated field is presented in Sec. IV.1, before we describe the experimental procedure and results for this two-qubit entangling gate in Secs. IV.2 and IV.3, respectively. We conclude this article with a discussion of phase-modulated vs. amplitude-modulated driving fields for entangling gates as well as a summary of results and an outlook on future work in Sec. V.
@@ II Experimental overview
We first investigate the double-dressed-state dynamics of a single qubit interacting with the c.m. mode of a two-ion Coulomb crystal. Then, we implement conditional quantum dynamics with two qubits. For all experiments, we use 171Yb+ ions trapped in a macroscopic linear Paul trap with radial and axial c.m. angular frequencies of $2\pi\times 380$ kHz and $\nu=2\pi\times 98.08$ kHz, respectively. The experimental setup is similar to Ref. [18], and it is only briefly described here. The two qubits are encoded in hyperfine states of the electronic ground state of 171Yb+ ions, $|0\rangle\equiv{|^{2}\rm S_{1/2},\rm F=0,\rm m_{\rm F}=0\rangle}$ and $|1\rangle\equiv{|^{2}\rm S_{1/2},\rm F=1,\rm m_{\rm F}=-1\rangle}$.
A static magnetic field gradient of $19.09(1)$ T/m applied along the trap axis ($z$ axis) yields individual Zeeman shifts of the two qubits and results in individual angular resonance frequencies, $\omega_{0}^{(1)}$ and $\omega_{0}^{(2)}$, of the magnetic dipole transition between qubit states. Both frequencies are near $2\pi\times 12.6$ GHz and differ by $2\pi\times 4.342(1)$ MHz. Thus, qubits are individually coherently controlled with low crosstalk using global rf radiation [21].
The entangling gate proposed and implemented here takes advantage of the state-dependent force induced by the static magnetic field gradient. This state-dependent force is due to a differential energy shift of the two internal states that make up a qubit, thus coupling the qubit states to the vibrational states of the ion crystal [9, 11]. Thus, both in the single-ion experiments reported in Sec. III, and in the two-ion experiments in Sec. IV,the qubit states and the ions’ motional states are coupled via MAGIC.
Here, the coupling strength between qubit states and motional states, the latter acting as a quantum bus between qubits [1], is given by the effective Lamb-Dicke parameter, $\eta=g_{F}\mu_{B}(\partial_{z}B)\nu^{-3/2}/\sqrt{2Nm_{\text{Yb}}}$ which is proportional to the magnetic field gradient, $\partial_{z}B$ [9]; $g_{F}$ is the hyperfine Landé $g$-factor of ${}^{171}\text{Yb}^{+}$ (Appendix A); $m_{\text{Yb}}$ is the mass of ${}^{171}\text{Yb}$; $\mu_{B}$ is Bohr’s magneton; and $N$ is the number of ions in a linear Coulomb crystal. In the experiments presented here $\eta=0.0329$. Therefore, when using MAGIC, laser light is not required for implementing conditional gates with trapped ions.
The ions are cooled close to their motional ground state in two stages. Initial Doppler cooling is followed by rf sideband cooling of both vibrational modes, giving a mean phonon number of 0.6(5) in the c.m. mode measured with sideband thermometry [53]. The heating rate of one trapped ion in the current experimental setup is 0.19(3) phonons/ms for this mode. We use an arbitrary waveform generator (AWG) to generate the phase-modulated rf driving fields that double dress both qubits [45, 54, 44].
@@ III Double dressing of single qubits
The MAGIC scheme in a static gradient field requires a magnetic-field-sensitive qubit transition, making the qubits susceptible to decoherence because of magnetic-field fluctuations. We demonstrate experimentally that the double-dressing field acts as continuous dynamical decoupling and improves the coherence time significantly by suppressing the effect of addressing frequency fluctuations as well as amplitude fluctuations of the rf driving field. The idea of using double-dressed states to counter noise has been established in previous work on single qubits in nitrogen-vacancy centers in diamond [55, 44, 56]. Here, we extend double dressing to individual trapped ions in a two-ion Coulomb crystal, and we investigate the coupling of the qubit resonance to vibrational modes by the phase-modulated driving field.
Figure 1: Illustration of the emergence of double-dressed states and the consequent protection against frequency and amplitude fluctuations. Thin double arrows indicate resonance frequencies. Thick double arrows indicate Rabi frequencies. The bare qubit is driven near resonance, $\omega_{0}$, with a time-dependent detuning achieved by phase modulation according to Eq. (5) (left: bare). In the dressed-state basis (middle: dressed), the drive with Rabi frequency $\Omega_{1}$ translates into the energy gap $\Omega_{1}$. Relative frequency fluctuations $\delta\omega_{0}$ between the driving field and the qubit are suppressed in the dressed basis for $\delta\omega_{0}\ll\Omega_{1}$. The phase modulation transforms into an effective on-resonance drive in the dressed basis, creating double-dressed states (right: double-dressed). In the double-dressed frame, the effective second drive translates into the double-dressed-states energy gap. Amplitude fluctuations in the first drive $\delta\Omega_{1}$ are suppressed in the double-dressed basis for $\delta\Omega_{1}\ll\Omega_{2}$.
The double-dressed scheme is illustrated for one qubit in Fig. 1 and is described in Appendix A in detail. In brief, the ionic qubit in a harmonic oscillator potential is modeled as [9]
@@
Here, $\sigma_{z}$ is the the Pauli operator describing the qubit’s eigenstates. We consider the c.m. vibrational mode of two ions and denote $b$ as the ladder operator of the motional quanta in the effective harmonic trapping potential. In the mathematical description in this article we set $\hbar=1$. In all experiments reported in this article, we trap two ions (i.e., $N=2$) while the number of ions simultaneously addressed by a dedicated rf driving field changes. For the single-qubit experiments discussed in this section as well as for the two-qubit experiments in Sec. IV, undesired crosstalk between qubits is negligible [21].
We drive one ion near resonance with a phase-modulated field such that
@@
where $\Omega_{1}$ is the Rabi frequency of the rf drive, $\sigma_{x}$ is a Pauli operator, and $\Omega_{2}$ is a parameter quantifying the phase-modulation amplitude.
When the driving field is applied [Eq. (2)], we obtain the Hamiltonian
$H_{I}=\frac{\Omega_{1}}{2}S_{z}+\frac{\Omega_{2}}{2}S_{x}\cos(\Omega_{1}t)+\nu b^{\dagger}b-\frac{\eta\nu}{2}S_{x}\left(b+b^{\dagger}\right)$ (3)
in the rotating frame with respect to $\omega_{0}\sigma_{z}/2+\Omega_{2}\sigma_{z}/2\cos\left(\Omega_{1}t\right)$. Here, $S_{x}=-\sigma_{z},S_{y}=\sigma_{y},S_{z}=\sigma_{x}$ are the dressed-basis operators. Dressing suppresses fluctuations of the bare state energy gap $\delta\omega_{0}\ll\Omega_{1}$ caused by either frequency fluctuations of the driving field or by fluctuations of magnetically sensitive bare qubit states which in turn may be caused by electric noise fields.
Phase modulation of the dressing field, in addition, suppresses fluctuations in the amplitude of $\Omega_{1}$: In an appropriate rotating frame, phase modulation creates an effective on-resonance second drive in the dressed basis. This second drive dresses the qubit a second time and thus dynamically decouples amplitude noise fluctuations in the drive $\delta\Omega_{1}$, which would otherwise reduce the coherence time of the dressed qubit (Fig. 1). These fluctuations are suppressed as long as ${\delta\Omega_{1}\ll\Omega_{2}}$. Furthermore, undesired coupling of the qubit to the motional state, which leads to a decay in the Rabi oscillations, is suppressed as long as $\Omega_{2}\gg\eta\nu$ and $|\Omega_{1}-\nu|\ll\Omega_{2}\ll|\Omega_{1}+\nu|$ (Appendix A gives a detailed description).
@@ III.1 Extending the coherence time
Figure 2: Relaxation time $T_{2,\rho}$ of a single qubit as a function of the modulation depth $\Omega_{2}/\Omega_{1}$. The qubit is initialized in state $|0\rangle$, which is a superposition state in the dressed-state basis. The phase-modulated driving field is continuously applied, and the coherence time $T_{2,\rho}$ is extracted from an exponential decay of the Rabi fringe contrast. At a modulation depth $\Omega_{2}/\Omega_{1}\approx 0.75$, the relaxation time is enhanced by 3 orders of magnitude compared to the scenario without a second dressing field. The statistical error bars for most data points are smaller than the markers.
The impact of the phase-modulated dressing field on a qubit’s coherence time is shown in Fig. 2. We study the relaxation of a single qubit in a two-qubit crystal initialized in state $|0\rangle$, while the phase-modulated driving field given in Eq. (2) is continuously applied. Note that this is an equal superposition state in the dressed basis; therefore, by fitting the decay of the oscillations’ contrast to an exponential decay, we can extract the coherence time $T_{2,\rho}$ of the dressed state.
When no modulation is applied, we observe ordinary Rabi oscillations (not shown) with frequency $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, close to the trap frequency $\nu=2\pi\times 98.08\leavevmode\nobreak$kHz. At low modulation strength, the coherence time is mainly limited by the near-resonant coupling of the qubit to the c.m. motional mode. With increasing modulation amplitude, $\Omega_{2}/\Omega_{1}$, the coherence time of the dressed state, $T_{2,\rho}$, increases.
@@
For $\Omega_{2}/\Omega_{1}=0.75$, used for the entangling gate mechanism in this work, we observe 3-orders-of-magnitude improvement in $T_{2,\rho}$, up to 0.27(2) s. The slightly reduced coherence time for $\Omega_{2}/\Omega_{1}\approx 1$ is due to second-order fluctuations of $\delta\Omega_{2}$ [44] and to the quantum Stark shift induced by counterrotating terms.
III.2 Coherent coupling to motional degrees of freedom